What is the diagonal distance in a rectangle with the dimensions of 20’ x 8’?
Answer by checkley79(3341)

The diagonal would be the hypotenuse of a 20 by 8 right triangle.
20^2+8^2=c^2
400+64=c^2
464=c^2
c=sqrt464
c=21.54 ans. for the diagonal (hypotenuse).